OpenStudy (anonymous):

The Miami Chargers a high school football team completed their season with a mean of 32 points scored for all of their home games. The final scores for each of their home games is listed below.

OpenStudy (anonymous):

How many points did the Miami Chargers score in their first home game? A. 22 B. 24 C. 26 D. 28

OpenStudy (anonymous):

x + 46 + 42 + 26 + 30 + 24 ------------------------- = 32 solve for x 6

OpenStudy (anonymous):

I did that

OpenStudy (anonymous):

and what did you get?

OpenStudy (anonymous):

It isn't working for me

OpenStudy (anonymous):

With 22 i got 354

OpenStudy (anonymous):

really. i got one of the answer choices. x + 168 --------- = 32 6 now multiply both sides by 6 x + 168 = 192 x = 24

OpenStudy (anonymous):

I just found it out whoops.